OpenStudy (anonymous):

Franklin's hair grew one-twelfth inches in one week. If his hair continues to grow at this rate, about how many inches would his hair grow in one year, or fifty-two weeks? 1 4 8 12

OpenStudy (anonymous):

1/12 * 52 = 4 to nearest inch

OpenStudy (chrisplusian):

if it grew i/52 inch in one week you can model that by saying\[(1/12) *1=12\]

OpenStudy (chrisplusian):

then (1/12) *52weeks = 52/12 =your answer